OpenStudy (lilkg77):

what is the point-slope equation of the line with slope -10 that goes through the point (1,4)? A. y-4=10(x-1) B. y+4=10(x+1) C. y-4=-10(x+1) D. y-4=-10(x-1)

OpenStudy (kat_s123):

y-4=-10(x-1)

OpenStudy (mathmale):

@kat_s123 : Thank you for trying to be helpful. In the long run it'd be a lot more helpful to other OpenStudy users if you'd please NOT provide answers, but rather help the other person find his or her own answers.

OpenStudy (kat_s123):

Okay.. well the formula is y - y1 = m(x-x1). y1 and x1 represents any point that you are given. and m represents the slope. And then you substitute your given info in and you can create the equation.